# Execution Guided Line-by-Line Code Generation

### v1

@@ 3.3 Execution-Guided Inference Loop
Our inference procedure extends standard autoregressive token generation by explicitly incorporating dynamic execution feedback via CFG. Starting from an initial prompt $p_{\text{pre}}$ (Equation 3), we autoregressively sample tokens $w_{i}$ from the CFG-conditioned distribution $M_{\text{CFG}}(w_{i}\mid p_{\text{sol}},p_{\text{dyn}})$, progressively constructing the solution sequence $p_{\text{sol}}$.
At each token-generation step, we reuse the dynamic signal $p_{\text{signal}}$ (Equation 8), injecting it into $p_{\text{sol}}$ at index $i_{\text{dyn}}$ to form $p_{\text{dyn}}$ as described in Equation 9. $p_{\text{signal}}$ itself is regenerated only upon completing a new line.
$w_{\text{sol}}^{i}=\arg\max M_{\text{CFG}}(w_{i}\mid p_{\text{sol}},p_{\text{dyn}}),\quad\quad p_{\text{sol}}=[p_{\text{pre}},w_{\text{sol}}^{0},\cdots,w_{\text{sol}}^{n-1}]$ (11)
To systematically leverage execution-based guidance, given an input task $\tau=(p_{\text{inst}},T,f_{\text{name}})$, we perform inference over all parameter combinations of completion candidates $s$, completion horizon $d$, sampling temperature $t$, and instruction prompt types. For each such parameter set, we iteratively run inference with progressively increasing CFG strength values $\gamma$. We return the first correct solution identified; otherwise, inference continues across remaining parameter combinations.
Instruction (from $p_{0}$) Write a python function to find the first non-repeated character in a given string. assert first_non_repeating_character("aabc") == "c"
assert first_non_repeating_character("abcabc") == None
assert first_non_repeating_character("abc") == "a"
assert first_non_repeating_character("ababc") == "c"
Dynamic signal instruction $p_{\text{dyn-inst}}$ starts here at $i_{\text{dyn}}$
Below are execution traces from running the response function after appending several possible future continuations. These continuations represent plausible ways the function might continue from its current state. They are not necessarily full solutions - some may be partial, exploratory, or incomplete.
For each candidate continuation, multiple test cases (invocations) were executed to observe its behavior under different inputs. Each entry includes: - A candidate version of the function - A specific test case used for invocation - The resulting execution trace for that test case These dynamic signals can help you better understand how different plausible continuations behave at runtime, and guide you toward a more accurate solution. Execution feedback for a single test $|T|=1$ and $s=2$ candidates # Function: def first_non_repeating_character(s): char_count = {} for char in s: char_count[char] = char_count.get(char, 0) + 1 for char in s: if char_count[char] == 1: return char return None # Invocation:
first_non_repeating_character("aabc")# Execution Trace:
s = ’aabc’, char_count = {} char = ’a’ -> char_count = {’a’: 1} char = ’b’ -> count = 1 -> return ’b’ # Function: def first_non_repeating_character(s): char_count = {} for char in s: char_count[char] = char_count.get(char, 0) + 1 for char in s: if char_count[char] == 2: return char return None # Invocation:
first_non_repeating_character("aabc")
# Execution Trace:
s = ’aabc’, char_count = {} char = ’a’ -> char_count = {’a’: 1} char = ’a’ -> char_count = {’a’: 2} char = ’b’ -> char_count = {’a’: 2, ’b’: 1} char = ’c’ -> char_count = {’a’: 2, ’b’: 1, ’c’: 1} char = ’a’ -> count = 2 -> return ’a’ ### Response: ‘‘‘python def first_non_repeating_character(s): char_count = {} for char in s: char_count[char] = char_count.get(char, 0) + 1
@@ 4 Experiments
We conduct our experiments using two LLMs across different parameter scales: DeepSeek-Coder-1.3B guo2024deepseek , which is small enough to run locally on our machines (NVIDIA GeForce RTX 2080 Ti and RTX 3090 GPUs), and a large open-source model, DeepSeek-V3-0324 liu2024deepseek which we use through a cloud inference endpoint. We use Fireworks AI which was chosen based on two criteria: a modest cost, and the availability of a log probability output that is required to perform the CFG (Equation 10). The full code used for our experiments is provided in the supplementary materials.
@@
As explained in section 3.3, given a task, multiple hyperparameters are tried to create a diversity of candidates. The following hyper-parameter sets were used in our experiments: $s=3$, $t\in\{0.7,0.75,0.85,0.95,1.2,1.5\}$, $d\in\{2,3,6,8\}$, $\gamma\in\{0,0.5,1,3\}$. Additionally, we evaluate both $p_{0}$ prompts, see section 3 and appendix Appendix A.
@@
Our evaluations are performed on two widely-used code-generation benchmarks: MBPP austin2021program and HumanEval chen2021evaluating , which consist of Python programming tasks described in natural language, accompanied by test cases for evaluating the correctness of generated solutions (MBPP: 500 tasks, HumanEval: 164 tasks). Additionally, we utilize the MBPP-ET and HumanEval-ET datasets dong2025codescore , enhanced versions containing extended and more challenging test cases, to rigorously evaluate the generalization and robustness of the generated solutions. These benchmarks are extensively used in the literature due to their clarity, diversity, and practicality. In addition, we evaluate our method on the challenging competitive programming benchmark, CodeContests li2022competition , which is designed to assess advanced algorithmic problem-solving abilities. On CodeContests, evaluations were performed using the ExecEval evaluation framework. (khan2024xcodeeval, )
We evaluate model-generated solutions using the standard accuracy metric used in benchmarks, defined as the percentage of problems for which the generated solution passes all provided test cases, reflecting the correctness and practical usability of the generated code.
@@
We compare our EG-CFG method against several established state-of-the-art methods for code generation and debugging:
  • Baseline LLM Few-shot prompting using the few-shot template introduced in DeepSeek-Coder’s official evaluation guo2024deepseek .
  • MGDebugger shi2024code : An iterative refinement method that combines test-cases assertions based feedback from an external Python interpreter with execution traces as simulated by the LLM.
  • MapCoder islam2024mapcoder : Employs multi-agent interactions to generate and iteratively refine code.
  • QualityFlow hu2025qualityflow : Incorporates agentic workflows to iteratively enhance generated code through quality-focused interactions.
  • LPW lei2024planning : Utilizes a structured two-phase workflow involving plan creation, plan verification, and iterative code refinements guided by runtime execution feedback.
@@ 4.1 Results
On the MBPP benchmark Table 1, our EG-CFG method outperforms all other approaches . Notably, EG-CFG achieves new state-of-the-art results on both MBPP and MBPP-ET benchmarks with the DeepSeek-V3-0324 model, outperforming previous results obtained by large closed-source models such as GPT-4 and Claude-Sonnet-3.5.
On the HumanEval benchmark Table 2, EG-CFG achieves highly competitive results compared to existing SOTA methods. The method that outperforms it (98.2% vs. 96.95%) is LPW when using the GPT-4o model, and MapCoder also achieves 96.95%. Crucially, when evaluated on the more challenging HumanEval-ET benchmark, EG-CFG achieves another new state-of-the-art result.
On the CodeContests benchmark Table 3, our EG-CFG method achieves a new state-of-the-art overall accuracy using the DeepSeek-V3-0324 model. This significantly surpasses the DeepSeek-V3-0324 Baseline LLM and MapCoder. Furthermore, EG-CFG on DeepSeek-V3-0324 also outperforms previous results from GPT-4 models using methods like LPW, CodeSim, and MapCoder.
To further understand the failure cases of our method, we manually reviewed all the tasks in MBPP that were not solved by DeepSeek-V3-0324. We found that out of the total 17 tasks that the model did not pass, nine of them contain invalid unit tests, with some also having incorrect reference solutions. In these nine cases, our method generated correct code that was marked as failed due to flawed benchmark tests. Full details are provided in the supplementary material.
We note that across all tested baselines, the publicly available code was highly sensitive to the specific model and could not be readily applied to DeepSeek models. We invested substantial effort in debugging and adapting the code to ensure it produced meaningful results that represented each baseline method as favorably as possible. Other methods, such as QualityFlow, have not release their code, preventing us from evaluating them on the DeepSeek models. While LPW did release public implementation, we encountered substantial technical issues during the execution of their published code on DeepSeek-Coder 1.3b model, resulting in unusually low scores despite our best debugging efforts on that benchmark.
@@
Table 2: Performance on the HumanEval and HumanEval-ET benchmarks. Our proposed EG-CFG achieves a new state-of-the-art overall accuracy on HumanEval-ET.
@@
Table 3: Performance on the CodeContests benchmark. Our proposed EG-CFG achieves a new state-of-the-art overall accuracy on CodeContests. The runs on Deep-Seek-V3-0324 are by us, and all other results are quoted from the literature. The LPW results was obtained on a custom test-set; the published code was not compatible with evaluating on DeepSeek. ∗∗Reported by the MapCoder paper (islam2024mapcoder, ). ∗∗∗Reported by the CodeSim paper (islam2025codesim, ).
@@
DeepSeek-Seek-V3-0324 Baseline LLM 41.81 0.00
DeepSeek-Seek-V3-0324 EG-CFG (Ours) 58.18 28.13
DeepSeek-Seek-V3-0324 MapCoder (islam2024mapcoder, ) 50.30 14.59
@@
We report the mean time to solve each task under full parallelization. Unlike iterative refinement methods, which require sequential feedback loops, EG-CFG treats each inference run (across different hyper-parameters for the same task) as independent, enabling parallel execution. Runtimes include early termination, i.e., once a task is solved, remaining runs are terminated. As can be seen in Table 4, our method is more efficient than MGDebugger and competitive with MapCoder. When comparing with LPW, our method is more efficient with the smaller model but slower with the larger model.
@@
We performed an ablation study on the MBPP and MBPP-ET benchmarks to evaluate various components of our method. The results are reported in Table 5. When omitting the beam search of Equation 4, which creates multiple solutions instead of a single completion, the performance of the methods drops and becomes much closer to the baseline performance.
The role of CFG is evident from the second ablation, in which a value of $\gamma=1$ is used in Equation 10. In this case, there is a clear drop in performance, although results are still clearly above the baseline.
Finally, a similar drop in performance is observed when replacing the detailed execution trace used as part of the dynamic signal with a minimal execution trace that captures only the final values and types of all variables in scope after the last line of the solution function has completed execution.
@@ 5 Conclusions
The empirical results demonstrate that EG-CFG achieves new state-of-the-art results of 96.6% accuracy on MBPP, 73.0% on MBPP-ET, 87.19% on HumanEval-ET and 58.18% on CodeContests using DeepSeek-V3-0324, surpassing both open-source and proprietary model-based approaches. The superior performance on both MBPP-ET and HumanEval-ET underscores the method’s robustness and effectiveness in generating reliable and accurate code under complex and extended test scenarios.
Notably, our approach demonstrates robust performance even with smaller models, achieving 83.2% accuracy using DeepSeek-Coder-1.3B, comparable to results from substantially larger models. This scalability highlights the effectiveness of execution signals as a guiding mechanism regardless of model capacity.
@@
Looking forward, this work opens several promising research directions. The execution-guided framework could be extended to more complex programming tasks requiring longer-horizon planning or multi-file interactions.
Additionally, the underlying principles of EG-CFG dynamically incorporating external semantic signals into generative processes could benefit other domains where generation benefits from grounding in external systems, such as database querying, formal verification, or simulation-based generation.
@@ 6 Limitations
First, the approach introduces computational overhead compared to standard inference methods. The beam search exploration, execution of multiple candidate continuations, and dual-distribution interpolation in the CFG mechanism collectively increase inference time. Though our parallel execution strategy mitigates this overhead somewhat, future work should explore more efficient methods for extracting and incorporating execution signals.
Second, EG-CFG’s effectiveness is contingent upon the availability of executable test cases that adequately exercise the target functionality. In real-world programming scenarios, comprehensive test cases may not always be available or easily generated by LLMs, potentially limiting the approach’s applicability.
Finally, because our inference loop is bottom-up, it does not exploit task decomposition, a strategy that has been shown to improve code generation shi2024code . Future work could integrate our sampling strategy with iterative refinement methods, task-decomposition methods or with top-down problem-inspection techniques shalev2024artificial to achieve even better performance.
Despite these limitations, EG-CFG represents a significant advancement in execution-aware code generation and provides a solid foundation for future research addressing these challenges.
@@ Appendix A The two $p_{\text{inst}}$ prompts used by our method
The two basic instruction prompts that are used by our method are provided in Figure 5 and Figure 6. The former is the standard DeepSeek-Coder 3-shot prompt guo2024deepseek , consisting of concise instructions accompanied by short illustrative examples, and the latter is an alternative prompt explicitly designed to encourage the generation of step-by-step solutions with more atomic logic.

### v2

@@ 3.3 Execution-Guided Inference Loop
Our inference procedure extends standard autoregressive token generation by explicitly incorporating dynamic execution feedback via CFG. Starting from an initial prompt $p_{\text{pre-sol}}$ (Equation 5), we autoregressively sample tokens $w_{i}$ from the CFG-conditioned distribution $M_{\text{CFG}}(w_{i}\mid p_{\text{sol}},p_{\text{dyn}})$, progressively constructing the solution sequence $p_{\text{sol}}$. At each token-generation step, we reuse the dynamic signal $p_{\text{signal}}$ (Equation 11), injecting it into $p_{\text{sol}}$ at index $i_{\text{dyn}}$ to form $p_{\text{dyn}}$ as described in Equation 12. $p_{\text{signal}}$ itself is regenerated only upon completing a new line.
$w_{\text{sol}}^{i}=\arg\max M_{\text{CFG}}(w_{i}\mid p_{\text{sol}},p_{\text{dyn}}),\quad\quad p_{\text{sol}}=[{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}p_{\text{pre-sol}}},w_{\text{sol}}^{0},\cdots,w_{\text{sol}}^{n-1}]$ (14)
3.4 Parallel Multi-Agent Execution
Given an input task as defined in Equation 1, we launch a parallel multi-agent inference process where each agent is assigned a unique configuration: candidate count $s$, generation horizon $d$, sampling temperature $t$, instruction prompt template $p_{0}$ and guidance strength $\gamma$. Once any agent finds a correct solution, it is immediately returned and all remaining agents are terminated. The full pseudo-code for both the execution-guided inference loop and the multi-agent controller is provided in Appendix C.
@@ 4 Experiments
We conduct our experiments using two LLMs across different parameter scales: DeepSeek-Coder-1.3B guo2024deepseek , which is small enough to run locally on our machines (NVIDIA GeForce RTX 2080 Ti and RTX 3090 GPUs), and a large open-source model, DeepSeek-V3-0324 liu2024deepseek , which we use through a cloud inference endpoint. We use Fireworks AI which was chosen based on two criteria: a modest cost, and the availability of a log probability output that is required to perform the CFG (Equation 13). The full code used for our experiments is provided in the supplementary materials.
@@
As explained in section 3.4, our method launches multiple parallel agents for each task. Each agent is assigned a different hyperparameter configuration. The following hyper-parameter sets were used in our experiments: $s=3$, $t\in\{0.7,0.75,0.85,0.95,1.2,1.5\}$, $d\in\{2,3,6,8\}$, $\gamma\in\{0,0.5,1,3\}$. Additionally, we evaluate both $p_{0}$ prompt templates, see section 3 and appendix Appendix A.
@@
Our evaluations use widely-adopted benchmarks: MBPP austin2021program (500 tasks) and HumanEval chen2021evaluating (164 tasks), along with their extended test versions MBPP-ET and HumanEval-ET dong2025codescore . To assess performance on more challenging tasks, we also evaluate on the DS-1000 data science benchmark Lai2022DS1000 (1000 tasks) and the CodeContests competitive programming benchmark li2022competition (using the ExecEval framework khan2024xcodeeval ). We report accuracy: the percentage of problems passing all test cases. To rigorously test generalization and prevent overfitting to public tests, evaluations on HumanEval, HumanEval-ET, MBPP-ET, CodeContests, and DS-1000 rely on hidden test cases inaccessible during inference. DS-1000’s structure, providing only a single input-output example per problem, further tests model robustness against reliance on examples.
@@
We compare our EG-CFG method against several established state-of-the-art methods for code generation and debugging: a Baseline LLM using the few-shot template from DeepSeek-Coder’s evaluation guo2024deepseek ; MGDebugger shi2024code , an iterative refinement method combining test-case feedback with LLM-simulated execution traces; MapCoder islam2024mapcoder , which employs multi-agent interactions; QualityFlow hu2025qualityflow , which incorporates agentic workflows for iterative enhancement; LPW lei2024planning , which utilizes a structured two-phase workflow with runtime execution feedback; and CodeSim islam2025codesim , a search-based method that adapts semantically similar code snippets.
@@ 4.1 Results
Using the DeepSeek-V3-0324 model, EG-CFG achieves new state-of-the-art (SOTA) results across all evaluated benchmarks. On MBPP and MBPP-ET (Table 1), it surpasses prior methods using large closed-source models like GPT-4 and Claude 3.5. On HumanEval (Table 2), it matches the state-of-the-art (LDB) at 99.4% and sets a new state-of-the-art on the challenging HumanEval-ET variant. Furthermore, it establishes new state-of-the-art on CodeContests (Table 3), significantly exceeding its own baseline, MapCoder, and previous GPT-4 based approaches like LPW and CodeSim, and also sets a new state-of-the-art on DS-1000 (69.9%, Table 4), outperforming CONLINE/GPT-4.
We note that across all tested baselines, the publicly available code was highly sensitive to the specific model and could not be readily applied to DeepSeek models. We invested substantial effort in debugging and adapting the code to ensure it produced meaningful results that represented each baseline method as favorably as possible. Other methods, such as QualityFlow, have not released their code, preventing us from evaluating them on the DeepSeek models. While LPW did release a public implementation, we encountered substantial technical issues during the execution of their published code on the DeepSeek-Coder 1.3b model, resulting in unusually low scores despite our best debugging efforts on that benchmark.
@@
Table 2: Performance on the HumanEval and HumanEval-ET benchmarks. Our proposed EG-CFG matches the state-of-the-art on HumanEval and achieves a new state-of-the-art on HumanEval-ET.
@@
LLaMA 3 LDB (zhong2024debug, ) 99.4 - - -
Claude-Sonnet-3.5 QualityFlow (hu2025qualityflow, ) 98.8 - - -
@@
Table 3: Performance on the CodeContests benchmark. Our proposed EG-CFG achieves a new state-of-the-art overall accuracy on CodeContests. The runs on DeepSeek-V3-0324 are by us, and all other results are quoted from the literature. The LPW results were obtained on a custom test-set; the published code was not compatible with evaluating on DeepSeek. ∗∗Reported by the MapCoder paper (islam2024mapcoder, ). ∗∗∗Reported by the CodeSim paper (islam2025codesim, ).
@@
DeepSeek-V3-0324 Baseline LLM 41.81 0.00
DeepSeek-V3-0324 EG-CFG (Ours) 60.6 32.29
DeepSeek-V3-0324 MapCoder (islam2024mapcoder, ) 50.30 14.59
DeepSeek-V3-0324 CodeSim (islam2025codesim, ) 52.72 18.76
@@
A key advantage of EG-CFG is its native parallelism, a design choice that aligns with the growing trend of leveraging large-scale, parallel compute in modern agentic systems. This architectural choice is a key differentiator from iterative refinement methods, which are inherently sequential and therefore cannot leverage parallel compute to accelerate work on a single task. In contrast, EG-CFG is designed to explore diverse reasoning paths simultaneously across multiple agents (as described in section 3.4). This fundamental architectural difference makes wall-clock time the most relevant and fair metric for comparison. As can be seen in Table 5, on the MBPP benchmark, our method is more efficient than MGDebugger and competitive with MapCoder. When comparing with LPW, our method is more efficient with the smaller model but slower with the larger model.
Table 4: Performance on the DS-1000 benchmark. Our proposed EG-CFG achieves a new state-of-the-art overall accuracy on DS-1000. Reported by the SelfEvolve paper (jiang2023selfevolve, ). ∗∗Reported by DS-1000 (Lai2022DS1000, ) official leaderboard.
Model Method Accuracy (%) RSR (%)
DeepSeek-V3-0324 EG-CFG (Ours) 69.9 50.73
DeepSeek-V3-0324 Baseline LLM 38.9 0.00
GPT-4 CONLINE (he2024conline, ) 68.0 -
GPT-4 Baseline LLM 60.2 -
GPT-3.5 Turbo SelfEvolve (jiang2023selfevolve, ) 57.1 -
GPT-3.5 Turbo Self Debugging (chen2023teaching, ) 53.0 -
GPT-3.5 Turbo DocPrompting (zhou2022docprompting, ) 45.50 -
GPT-4o Baseline LLM∗∗ 59.9 -
Claude 3.5 Sonnet Baseline LLM∗∗ 54.3 -
DeepSeek-Coder-V2-SFT Baseline LLM∗∗ 53.2 -
Qwen2-72B-Instruct Baseline LLM∗∗ 52.8 -
Fairness of Comparison
To demonstrate EG-CFG’s qualitative advantage is not just a larger compute budget, we drastically increased baseline token usage. On MBPP (DeepSeek-V3-0324), we raised MGDebugger and MapCoder retry counts 40-fold (from 5 to 200). Despite this 40x compute increase, gains were marginal : MGDebugger improved by only 6.8 percentage points, from 86.8% to 93.6% (solving 34 of its 86 prior failures). MapCoder’s improvement was even smaller, at 1.6 percentage points, from 87.2% to 88.8% (solving 8 of its 64 prior failures). Both remained significantly below EG-CFG’s 96.6% accuracy , showing our performance gap is due to the execution-guided feedback loop’s quality, rather than the amount of computation.
@@
We performed an ablation study on the MBPP and MBPP-ET benchmarks to evaluate various components of our method. The results are reported in Table 6. When omitting the beam search of Equation 6, which creates multiple solutions instead of a single completion, the performance of the method drops and becomes much closer to the baseline performance. The role of CFG is evident from the second ablation, in which a value of $\gamma=1$ is used in Equation 13. In this case, there is a clear drop in performance, although results are still clearly above the baseline. Finally, a similar drop in performance is observed when replacing the detailed execution trace used as part of the dynamic signal with a minimal execution trace, formally defined as the final event of the execution trace (Equation 10).
@@ 5 Conclusions
The empirical results demonstrate that EG-CFG achieves new state-of-the-art results across a diverse set of benchmarks. Using DeepSeek-V3-0324, it achieves 96.6% accuracy on MBPP, 73.0% on MBPP-ET, 99.4% on HumanEval, 89.02% on HumanEval-ET, 69.9% on DS-1000, and 60.6% on CodeContests, surpassing both open-source and proprietary model-based approaches. The superior performance on both MBPP-ET and HumanEval-ET underscores the method’s robustness and effectiveness in generating reliable and accurate code under complex and extended test scenarios. Notably, our approach demonstrates robust performance even with smaller models, achieving 83.2% accuracy using DeepSeek-Coder-1.3B, comparable to results from substantially larger models like GPT-4. This scalability highlights the effectiveness of execution signals as a guiding mechanism regardless of model capacity.
@@
Looking forward, this work opens several promising research directions. The execution-guided framework could be extended to more complex programming tasks requiring longer-horizon planning or multi-file interactions. Additionally, the principles of EG-CFG, which dynamically incorporate external semantic signals into generation, could benefit domains that rely on grounding in systems, such as database querying, formal verification, or simulation-based tasks. More broadly, EG-CFG represents a shift in generative modeling beyond static, pattern-based generation toward responsive, context-aware methods informed by environmental interaction. As models scale and applications expand, such approaches can drive the development of systems that generate high-quality outputs while reasoning about their real-world behavior, enabling more reliable and aligned generation.
@@ 6 Limitations
First, the approach introduces computational overhead compared to standard inference methods. The beam search exploration, execution of multiple candidate continuations, and dual-distribution interpolation in the CFG mechanism collectively increase inference time. Though our parallel execution strategy mitigates this overhead somewhat, future work should explore more efficient methods for extracting and incorporating execution signals. Second, EG-CFG’s effectiveness is contingent upon the availability of executable test cases that adequately exercise the target functionality. In real-world programming scenarios, comprehensive test cases may not always be available or easily generated by LLMs, potentially limiting the approach’s applicability.
Finally, because our inference loop is bottom-up, it does not exploit task decomposition, a strategy that has been shown to improve code generation shi2024code . Future work could integrate our sampling strategy with iterative refinement methods, task-decomposition methods or with top-down problem-inspection techniques shalev2024artificial to achieve even better performance. Despite these limitations, EG-CFG represents a significant advancement in execution-aware code generation and provides a solid foundation for future research addressing these challenges.
@@ Appendix A The two $p_{\text{inst}}$ prompts used by our method
The two basic instruction prompts used by our method are provided in Figure 5 and Figure 6. The former is the standard DeepSeek-Coder 3-shot prompt guo2024deepseek , consisting of concise instructions accompanied by short illustrative examples, and the latter is an alternative prompt explicitly designed to encourage the generation of step-by-step solutions with more atomic logic.
@@
Appendix B Example Dynamic Prompt Structure
This appendix provides a concrete example illustrating the structure of the dynamic prompt $p_{\text{dyn}}$ used by our EG-CFG method (described at the end of Section 3.1). Figure 7 below shows how the different components are assembled into the dynamic prompt $p_{\text{dyn}}$. It displays the original task instruction ($p_{\text{inst}}$), followed by the dynamically injected signal ($p_{\text{signal}}$). This signal begins with the dynamic signal instruction ($p_{\text{dyn-inst}}$) and includes execution traces for each candidate, obtained in this example by executing $s=2$ candidates against a single test case ($|T|=1$). Finally, the figure shows the model’s partially completed response generated up to that point.
Instruction from $p_{\text{inst}}$: Write a python function to find the first non-repeated character in a given string.
assert first_non_repeating_character("aabc") == "b"
assert first_non_repeating_character("abcabc") == None
assert first_non_repeating_character("abc") == "a"
assert first_non_repeating_character("ababc") == "c"
Dynamic Signal Instruction ($p_{\text{dyn-inst}}$) starts at $i_{\text{dyn}}$
Below are execution traces from running the response function after appending several possible future continuations. These continuations represent plausible ways the function might continue from its current state. They are not necessarily full solutions - some may be partial, exploratory, or incomplete. For each candidate continuation, multiple test cases (invocations) were executed to observe its behavior under different inputs. Each entry includes: - A candidate version of the function - A specific test case used for invocation - The resulting execution trace for that test case These dynamic signals can help you better understand how different plausible continuations behave at runtime, and guide you toward a more accurate solution.
  Execution feedback for a single test ($|T|=1$) and $s=2$ candidates:  # Function: def first_non_repeating_character(s): char_count = {} for char in s: char_count[char] = char_count.get(char, 0) + 1 for char in s: if char_count[char] == 1: return char return None # Invocation: first_non_repeating_character("aabc")
# Execution Trace:
s = ’aabc’, char_count = {} ... char = ’c’ -> char_count = {’a’: 2, ’b’: 1, ’c’: 1} char = ’b’ -> count = 1 -> return ’b’   # Function: def first_non_repeating_character(s): char_count = {} for char in s: char_count[char] = char_count.get(char, 0) + 1 for char in s: if char_count[char] == 2: return char return None # Invocation: first_non_repeating_character("aabc")
# Execution Trace:
s = ’aabc’, char_count = {} ... char = ’c’ -> char_count = {’a’: 2, ’b’: 1, ’c’: 1} char = ’a’ -> count = 2 -> return ’a’   ### Response: ‘‘‘python def first_non_repeating_character(s): char_count = {} for char in s: char_count[char] = char_count.get(char, 0) + 1
Appendix C EG-CFG Pseudo-Code
To complement the formal description in Section 3, we provide full pseudo-code for our method in this appendix. Algorithm 1 outlines the controller responsible for coordinating a multi-agent parallel inference process, where each agent explores diverse reasoning paths using a unique configuration. Each agent independently invokes the core decoding routine defined in Algorithm 2, which integrates dynamic execution feedback via classifier-free guidance (CFG). This loop incrementally generates a solution by incorporating real-time execution feedback into the prompt.
Algorithm 1 Multi-Agent Parallel Inference Controller
1:Task $\tau=(p_{\text{task}},T,f_{\text{name}})$ (Eq.1), Model $M$, EG-CFG Configurations $\mathcal{H}=\{(p_{0},s,d,t,\gamma)\}$
2:Launch parallel agents $A=\{\text{a}_{h}:h\in\mathcal{H}\}$
3:for all agents $a_{h}\in A$ with config $(p_{0},s,d,t,\gamma)$ in parallel do
4:  ${p}_{\text{inst}}\leftarrow\textsc{BuildInstructionPrompt}(p_{0},\tau)$ $\triangleright$ Build instruction prompt, Equation 3
5:  $p_{\text{sol}}\leftarrow\textsc{EG-CFG-InferenceLoop}(p_{\text{inst}},T,M,s,d,t,\gamma)$
6:  if $\mathrm{Execute}(p_{\text{sol}},t_{j})=\text{success},\forall t_{j}\in T$ then $\triangleright$ Verify solution, Equation 2
7:    return $p_{\text{sol}}$ $\triangleright$ Return first valid solution and terminate other agents
8:  end if
9:end for
10:return null $\triangleright$ Failure: No correct solution found
Algorithm 2 Execution-Guided Classifier-Free Guidance (EG-CFG) Inference Loop
[…]
3:Locate $i_{\text{solution}}$, $i_{\text{dyn}}$ in $p_{\text{pre}}$
[…]
7:while true do
[…]
9:    $\mathcal{C}_{raw}\leftarrow[]$ $\triangleright$ Initialize list for executable candidates
10:    for all each $j=1,\dots,s$ do
11:     Initialize $c_{j}\leftarrow[]$
12:     while $\mathrm{CountLines}(c_{j})<d$ do $\triangleright$ Generate a candidate continuation, Equation 6
13:      Sample $w^{k}_{c_{j}}\sim M(w_{k}\mid p_{\text{sol}},w^{<k}_{c_{j}};t)$
14:      Append $w^{k}_{c_{j}}$ to $c_{j}$
15:     end while
16:     $\hat{c}^{j}\leftarrow\mathrm{ExtractExecutable}(c^{j})$ $\triangleright$ Extract executable part, Equation 7
17:     Append $\hat{c}^{j}$ to $\mathcal{C}_{raw}$
18:    end for
19:    $C\leftarrow\mathrm{Unique}(\mathcal{C}_{raw})$ $\triangleright$ Filter for unique candidate continuations, Equation 8
20:    $\mathcal{E}\leftarrow[]$ $\triangleright$ Initialize list for feedback tuples
21:    $\triangleright$ Candidate execution against test cases are running in parallel, Equation 9
22:    for all each $\hat{c}^{j}\in C$ do
23:     for all each $t_{m}\in T$ do
24:      $e^{j,m}\leftarrow\mathrm{ExtractExecutionFeedback}(\hat{c}^{j},t_{m})$
25:      Append $(\hat{c}^{j},t_{m},e^{j,m})$ to $\mathcal{E}$
26:     end for
27:    end for
28:    $p_{\text{signal}}\leftarrow[p_{\text{dyn-inst}},\mathcal{E}]$ $\triangleright$ Construct the full dynamic signal, Equation 11
29:  end if
30:  $p_{\text{dyn}}\leftarrow[\,p_{\text{sol}}[:i_{\text{dyn}}],\,p_{\text{signal}},\,p_{\text{sol}}[i_{\text{dyn}}:]\,]$ $\triangleright$ Inject signal into the prompt, Equation 12
31:  $w_{i}\leftarrow\arg\max M_{\text{CFG}}(w\mid p_{\text{sol}},p_{\text{dyn}})$ $\triangleright$ Sample next token using CFG, Eq.13 Eq.14
32:  Append $w_{i}$ to $p_{\text{sol}}$ $\triangleright$ Append new token to the solution, Eq.13 Eq.14
33:  if $w_{i}=\texttt{EOT}$ or $\text{length}(p_{\text{sol}})\geq N_{\text{max}}$ then
34:    return $p_{\text{sol}}$
35:  end if
36:end while